OpenStudy (anonymous):

what value of b will make the trinomial 9x^2 + bx + 1/4 a perfect square?

OpenStudy (anonymous):

\[9x^2 + bx + \frac 14 = (3x)^2 + 2.3x.\frac 12 + \left( \frac 12 \right)^2 = \left( 3x+\frac 12 \right)^2 \] So , \(b = 3\)